1) Assume that a company expects to produce 10,900, 11,900, and 13,900 units of finished goods in January, February, and March, respectively. Each unit of finished goods requires 3 pounds of raw material and each pound of raw material costs $3.25. The company always maintains an ending raw materials inventory equal to 20% of next month’s production needs. What is the amount of expected raw materials purchases for February?

Multiple Choice

  • $119,925

  • $83,025

  • $116,325

  • $71,745

2) Assume that June’s production budget showed required production of 431,000 units, desired ending finished goods inventory of 22,000 units, and beginning finished goods inventory 9,000 units. What were June’s budgeted unit sales?

Multiple Choice

  • 444,000 units

  • 418,000 units

  • 412,500 units

  • 456,500 units

3)Assume that the amount of one of a company’s variable expenses in its flexible budget is $40,000. The actual amount of the expense is $42,000 and the amount in the company’s planning budget is $43,000. The activity variance for this expense is:

Multiple Choice

  • $1,500 U.

  • $1,500 F.

  • $3,000 F.

  • $3,000 U.

4)Assume that the cost formula for one of a company’s variable expenses is $5.00 per unit. The company’s planned level of activity was 2,000 units and its actual level of activity was 2,200 units. The actual amount of this expense was $10,160. The spending variance for this expense is:

Multiple Choice

  • $2,280 F.

  • $1,440 F.

  • $840 F.

  • $2,280 U.

5)Assume that the cost formula for one of a company’s variable expenses is $5.00 per unit. The company’s planned level of activity was 2,000 units and its actual level of activity was 2,200 units. The spending variance was $950 unfavorable. What is the actual amount of this expense?

Multiple Choice

  • $8,950

  • $9,950

  • $11,000

  • $11,950

6)Assume that a company’s planned level of activity was 2,000 units and its actual level of activity was 2,200 units. The spending variance for one of its fixed expenses was $200 favorable. The actual amount of the fixed expense was $10,600. What amount of this expense would be included in the company’s planning budget?

Multiple Choice

  • $8,800

  • $11,800

  • $9,800

  • $10,800
Answer #1

here the solution for the MCQs are as below

1) here it is given company can purchase current month requirement plus 20 % of next monts requirement so fo the month of Feb. 20% of 11900  is already purchased in the month of Jan. so now it need to purchase only 80% of 11900+20% of 13900(i.e., march ) total is 9520+2780. each unit requires 3 pounds of materials AND EACH POUND COSTING TO $3.25 THEREFORE TOTAL PURCHASE AMOUNT IS (9520+2780)*3*3.25= $119925 OPTION A

2) here sales volume = budgeted production + opening inventory - desired inventory

431000+9000-22000=418000 units

3). 3000 F (favorable) because here as per flaxible budget actual variance is less than planned budget cost of $3000

4). here variable cost unique feature is its per unit cost is constant hence actual cost incurred to produce 2200 units is 2200*5=11000 but actual is only 10160 i.e., actual is less than standard by 840 favorable

11000-10160= 840 F

5). here given per unit variable cost is $5 and actual production is 2200 units so standard cost is 2200units*$5= 11000 but here given is $950 unfavorable means more than standard therefore toatl actual cost is; standard cost +unfavorable expenses i.e., 11000+950= $11950.

6). as i mentioned in earlier in the above solved questions favorable mens actual cost lessthan standard so here favorable cost is $200 and actual cost is 10600 so, total standard cost included in budget for planning is sum of both

i.e., 10600+200= $10800.

thank you !!!